OpenStudy (anonymous):

So I have two questions on some make up work, which is about proofs by mathematical induction. Use Mathematical induction to prove that the statement is true for all natural numbers n 1. 5+8+11+...(3n+2)=n(3n+7)/2 2. 1+3+5+...+(2n-1)=n^2

OpenStudy (fibonaccichick666):

Ok, so what are the steps for induction.

OpenStudy (anonymous):

First you prove it for n=1, then you prove for k=n, and then for k+1 I think?

OpenStudy (fibonaccichick666):

well I wouldn't say prove, you just do a base case where n=1

OpenStudy (fibonaccichick666):

then you assume it is true for every n

OpenStudy (fibonaccichick666):

finally you have to show that it works for n+1 using what you have assumed

OpenStudy (fibonaccichick666):

So let's try the first one. What is the base case?

OpenStudy (anonymous):

the base case is (3n+2)= n(3n+7)/2? I'm not completely sure

OpenStudy (fibonaccichick666):

Base case is just say n=1 and see if the right side works, so I like to structure mine like this Base case: Let n=1 3n+2=3(1)+2=5 \(\huge{√}\) \(\frac{n(3n+7)}{2}=\frac{(1)(3(1)+7)}{2}=\frac{10}{2}=5\) \(\huge{√}\)

OpenStudy (fibonaccichick666):

See how they check out? All you are doing is testing the hypothesis

OpenStudy (anonymous):

ok so one side has to equal the other, so the two bits in the parentheses have to balance out

OpenStudy (fibonaccichick666):

uh yea I think you are on the right page

OpenStudy (fibonaccichick666):

So now, you write: Inductive step: Assume this is true for every n. Must show it is true for n+1.

OpenStudy (anonymous):

So do you just substitute n+1 into the parentheses where the n is?

OpenStudy (fibonaccichick666):

well sort of, you know how we just assumed that the hypothesis was correct for all n so now we have that mumbo jumbo+ (3(n+1)+2) and we want to finangle that into if we put n+1 instead of n into that mumbo jumbo hypothesis

OpenStudy (fibonaccichick666):

sorry, I'm a lazy typer lol

OpenStudy (anonymous):

okay so if we solve it out we are supposed to make both sides equal again? and no it's cool I understand :)

OpenStudy (fibonaccichick666):

yea essentially

OpenStudy (fibonaccichick666):

you just do some fun algebra to make them equal again, then inductively, this proves it for every n>1

OpenStudy (anonymous):

okay, thank you so much! I will try that. Hopefully it will work out fine.

OpenStudy (fibonaccichick666):

yea, if you want post your steps here whether it be in a picture of your work or typed and I'll check it, just tag me if you have any more problems

OpenStudy (anonymous):

Thanks! It's really nice of you to help me. Have a good night (or day)

OpenStudy (fibonaccichick666):

haha thanks, I'll be up for a while I have a final at 9a and need to write the essay for that time still :X but don't sweat this. Inductive proofs were the first that I learned because they have a set step by step method, eventually it just clicks. You'll get the hang of it soon

OpenStudy (anonymous):

That doesn't sound too great ,so get some rest. All the same, best of luck on your final and essay! I think I might be getting it though, it's just a simple manner of balancing each side multiple times and showing why it works.

OpenStudy (fibonaccichick666):

Thanks! and yea, just make sure you don't set them equal or if you do, you only manipulate the one side

OpenStudy (fibonaccichick666):

It's just like proving the trig identities

OpenStudy (anonymous):

so far the steps my notes have say this: 1. Show the formula works for n=1 2. Assume the statement is true for natural number k 3. Prove statement is true for k+1 I realize this different from what I guessed earlier, so I though I should post this. And trig identities were not my strong suit, I'm afraid.

OpenStudy (fibonaccichick666):

ah, and some people use k others use n. I personally like n

OpenStudy (anonymous):

Yes, that's what I have, but honestly I think both work

OpenStudy (fibonaccichick666):

but yea(trig idents were not my stong suit either lol) , just follow the steps. Memorize the format and the proofs will follow

OpenStudy (anonymous):

That's what I was having issues with, but now it's a lot clearer.

OpenStudy (fibonaccichick666):

It's a dummy variable, doen't matter, could use gamma or ro or smiley face for all it matters

OpenStudy (anonymous):

Well, I have to go and finish it off, so best of luck on your essay and final! And perhaps I will cover my next test in smiley faces, lol.

OpenStudy (fibonaccichick666):

Thanks lol and feel free just make sure you define it properly :P

OpenStudy (anonymous):

I will. Goodnight! :P